OpenStudy (anonymous):

graph the equation by plotting 3 points -y=x+1

OpenStudy (anonymous):

plug any value in for x. for example, x=-5, 0, and 3 -y=-5+1, -y=-4. so the ordered pair would be (-5, 4) -y=0+1, -y=1. this ordered pair would be (0, -1) -y=3+1, -y=4. this ordered pair would be (3, -4) just plot those three points and connect them

OpenStudy (anonymous):

THANK YOU THE THIS OTHER GIRL MADE ME GET IT WRONG CAN YOU HELP ME WITH ANOTHER PLEASE =)???